OpenStudy (anonymous):

The value of y varies directly with the value of x, and y = –8 when x = 4. Which equations represent the relationship between y and x? Choose exactly two answers that are correct. A. B. C. D.

OpenStudy (anonymous):

@warsop87

OpenStudy (anonymous):

@majesty69

OpenStudy (majesty69):

what are the options?

OpenStudy (anonymous):

A y/x=1/2 B. y/x=-2 C. -2x=y D. x=2y

OpenStudy (anonymous):

Actually no its not, Hang on let me double check

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

Perhaps if he comes he will give better advise then me

OpenStudy (anonymous):

okay thx